### v1

@@ 1 Introduction
High-precision predictions for particle colliders rely more than ever on the computation of multi-loop Feynman integrals. For this reason, the study of Feynman integrals and their analytic structure has become a major area of research in theoretical particle physics (see, e.g., [1, 2] for reviews). A central challenge in these computations is the reliable prediction of Feynman integral singularities as functions of kinematic parameters, without the need for a direct evaluation of the integral. In this paper, we tackle this problem by introducing [Uncaptioned image] (Singularities of Feynman Integrals Automatized), a Mathematica package designed for this purpose, which we release for public use:
SOFIA GitHub repository [3].
@@
The most common application is in computing Feynman integrals using the method of differential equations [4]. In the simplest cases, these equations take the form
@@
where $\vec{I}$ is a vector of basis Feynman integrals with $\mathrm{d}$ refers to the differential in the kinematic variables. On the right-hand side, $\mathbf{\Omega}_{i}$ are matrices with entries rational in the kinematics and the spacetime dimension $\mathrm{D}$. In fact, one often looks for bases in which all $\mathbf{\Omega}_{i}$ are linear in the dimensional-regularization parameter $\varepsilon=\frac{4-\mathrm{D}}{2}$ [5]. In this case, the resulting Feynman integrals $\vec{I}$ are often expressible in terms of multiple polylogarithms. The arguments of the logarithms, $W_{i}$, are known as the letters. Being able to predict these letters is currently one of the main bottlenecks in computations of polylogarithmic integrals: once $W_{i}$’s are known, one can fit the coefficients $\mathbf{\Omega}_{i}$ and consequently compute $\vec{I}$ using the differential equation together with some boundary conditions.
@@
Figure 1: Examples of cutting-edge diagrams and with timings (in minutes) to obtained singularities and letters. See [3] for expressions. Dashed lines denote massless particles and otherwise mass labels are displayed along the edges. The choice of SolverBound is determined by the maximum value for which the letters could be subsequently obtained. For diagram (c), although the singularity analysis finished in a reasonable time, the extraction of letters did not complete (hence the “?”) for any reasonable choice of SolverBound. For diagram (a), we used a machine with two Intel Xeon Gold CPUs and 512GB memory, while for (b) and (c) we used a MacBook Pro with an M3 chip and 16GB memory.
@@
Since problems involving analytic structure are notoriously subtle, let us be more precise about the type of questions that we aim to address. A singularity refers to a pole or a branch point in the (complexified) kinematic space. Let us collectively denote the kinematic variables as $\vec{s}\in\mathbb{C}^{\mathfrak{s}}$: they consist of the $\mathfrak{s}$ Mandelstam invariants and masses involved in the problem. While non-perturbatively scattering amplitudes might have a much richer singularity structure, in perturbation theory it is expected that singularities of the type
@@
where $f_{j}=f_{j}(\vec{s})$ is a polynomial and $A$, $B$ are rational numbers. For example, $\sim\sqrt{f_{j}}\log f_{j}$ is one possibility. Note that this refers to a local behavior of the amplitude near the locus $f_{j}=0$. It is much simpler than the global behavior which has to be compatible with (1.2) for a list of all possible $f_{j}$’s. Indeed, it is known that Feynman integrals are expressed in terms of complicated functions such as multi-polylogarithms, elliptic functions, periods of higher-genus Riemann surfaces and Calabi–Yau manifolds, etc. [6, 7] even though their local behavior (1.2) remains simple. The goal of singularity analysis is to exploit this simplicity.
@@
In the literature, the aforementioned singularities $f_{j}=0$ are also interchangeably known as Landau singularities or anomalous thresholds [8, 9, 10, 11]. The equation $f_{j}=0$ is a condition for “billiard ball scattering”: a network of classical on-shell particles following their classical, though possibly complexified, paths. The simplest case corresponds to positive energies and real angles [12], i.e., the physical region. But the interpretation extends to the rest of the physical sheet, and in fact any sheet, after relaxing positivity and reality conditions; see, e.g., [13].
@@
There has been an explosion of recent activity in the topic of determining computing singularities of Feynman integrals. Let us mention a couple of algorithms that have been implemented as parts of public packages and can handle state-of-the-art Feynman integrals. HyperInt implements a reduction algorithm that finds a superset of Landau singularities [14]. PLD.jl implements an algorithm based on toric geometry that finds a computable subset of Landau singularities [15, 16]. Both of these approaches work in the Schwinger-like parametrizations of Feynman integrals. A database comparing the two codes can be found in [17]. The package Baikovletter attempts to find symbol alphabets using Baikov representations [18]. We refer the readers to Sec. 2.2 for more detailed comparison and the introduction of [16] for a more complete list of references.
Importantly, it was recently observed that applying singularity analysis to the elastic unitarity constraint (relating the imaginary part of a diagram to its “cut” sub-diagrams) yields a powerful recursion formula for the calculation of Landau singularities in the loop momentum space [19]. This method enables efficient computation of the leading Landau singularities in two-particle reducible diagrams, successfully reproducing known results from the database [17] while also making new predictions, including all-loop examples.
@@
The first main novelty comes from a new variant of Baikov representations [20, 21] adapted to minimize the number of integration variables and polynomials involved in the singularity analysis. [Uncaptioned image] implements an optimization technique that results in a minimal integral representation (according to a metric explained in Sec. 2.1) without any user input besides the Feynman diagram under consideration. We believe it could come useful in other aspects of Feynman analysis that rely on Baikov representations, such as integration-by-parts reduction or analyzing Feynman geometries [22, 23, 24, 25, 26, 27, 28, 29, 30, 31]. We illustrate the latter application in Sec. 4.2. Likewise, we extend our analysis to eikonalized Feynman integrals, such as those in post-Minkowskian expansions with an example in Sec. 4.3.
@@
The second key innovation is the integration of advanced polynomial system solving methods for singularity analysis into the optimized Baikov representation. In particular, we implement a variant of the Fubini reduction algorithm [32] that results in an upper bound on the singularity locus. We also implement a Mathematica interface to PLD.jl, which allows one to seamlessly perform a numerical and symbolic analysis on a given diagram. We emphasize that this algorithm can be applied to any Euler integral and give a couple of examples of applications to cosmological and energy correlators in Sec. 4.4.
@@
The third and final significant novelty is the implementation of graph-theoretical tools to break down and speed up the computation of singularities. It has long been observed that the symmetries of a Feynman diagram lead to an identification between Schwinger parameters of the leading physical sheet singularity [33, 34] (see [35] for a modern application in the context of the S-matrix bootstrap). In this work, we extend this idea by relating different sets of subleading singularities through the construction and identification of subtopologies that are connected via symmetry transformations accounting for variations in their kinematics. This procedure effectively mitigates the factorial growth in the number of subtopologies that would otherwise need to be analyzed.
@@
Finally, we use the recent implementation of the publicly available packge Effortless [36] to provide an end-to-end function that takes a diagram as the input and provides a candidate symbol alphabet as the output. By “candidate” we mean that the alphabet might be overcomplete. An example is provided in Sec. 4.1.
@@
We believe that [Uncaptioned image] establishes substantial new milestones in the singularity analysis of Feynman integrals, producing results that were previously unattainable. The ancillary files [3] contain several new predictions for singularities and alphabets, including those presented in Fig. 1, which exemplify how [Uncaptioned image] can be applied to next-frontier examples, namely, $HZ+2$ jets, $H+$jet, and multi-jets production through light- (for (a)) or heavy-quark (for (b) and (c)) loop-induced gluon fusion at the LHC.
@@
We hope that [Uncaptioned image] serves as a tool to explore new avenues and sheds light on the analytic structures of perturbative scattering amplitudes that were previously inaccessible.
@@
We provide theoretical background in Sec. 2, including the discussion of different variants of the Baikov representation, singularity analysis of integrals, and converting from singularities to symbol alphabets. The manual of [Uncaptioned image] is given in Sec. 3 with an outline of all the functions and options. Sec. 4 focuses on examples and various applications. Finally, future directions are discussed in Sec. 5.
@@ 2.1 Baikov representations
There are a few ways to implement the change of variables (2.3). The general idea dates back to the S-matrix theory in the 1960s [37, 38, 39], although in the more recent literature on Feynman integrals parametrizations of this type are referred to as Baikov representations [20]. In the global Baikov representation, one transforms all variables at the same time. This results in the representation of the form
@@
The central object in the global Baikov representation is $\mathcal{G}(\vec{x},\vec{s})$ which arises as the general Jacobian for the change of variables. It can be expressed as a Gram determinant of a certain matrix. The integration contour $\Gamma^{\prime}$ is determined by $\vec{x}$ such that $\mathcal{G}(\vec{x},\vec{s})>0$. Likewise, $\tilde{\mathcal{G}}(\vec{s})$ is another Gram determinant that depends only on external kinematics and hence acts as an overall normalization. The derivation in App. A makes it clear that one can include different powers of propagators as well as non-trivial numerators with minimal modifications. For completeness, we review these aspects in App. A (see also [40, 41]).
The alternatives proposed by Frellesvig and Papadopoulos are loop-by-loop Baikov representations [22]. The idea is simply to Baikov parametrize each loop of the diagram sequentially in a given order. For each loop, we look for a shift in the loop momentum $\ell_{a}$ so that fewer invariants—such as $\ell_{a}\cdot p_{i}$ and $\ell_{a}\cdot\ell_{b}$—appear before changing variables from the $\ell$’s to the $x$’s. This procedure has the effect of decreasing the total number of integration variables $N\leqslant N^{\prime}$. Typically, the minimal/optimal $N$, is [42]
@@
Due to complicated integration contours, Baikov representations are typically not used in the explicit evaluation of Feynman integrals. However, they are extremely useful for integral reduction, integration-by-parts identities, singularity analysis, and other applications that rely on the simplicity of the integrand. The main advantage comes from the application of generalized unitarity. Generalized cuts can be taken by simply replacing the integration contour with residues around the relevant propagators (similarly, unitarity or Cutkosky cuts can be taken after an extra modification of the integration contour). For example, the maximal cut is obtained by
@@
in either global or loop-by-loop Baikov approach (i.e., $\mathcal{N}=N\,,N^{\prime}$). This replacement exists if the original integration domain, $\Gamma$ or $\Gamma^{\prime}$, restricted to $x_{1}=x_{2}=\ldots=x_{\mathcal{N}}=0$ is non-empty (more on that is discussed in Sec. 3.3). Note that after taking the cut, $\mathcal{N}-E$ variables still have to be integrated over.
@@
We implement a search algorithm that optimizes for the most efficient loop-by-loop Baikov representation without any user input. It proceeds in two steps. First, when possible, the loops are ordered automatically so that the number of external momenta $\mathcal{E}_{a}$ (associated with each loop momentum $\ell_{a}$, for $1\leqslant a\leqslant L$) increases from the smallest to the largest. [^1] This strategy is inspired by the recursive elastic unitarity approach of [19], where the selection and ordering of cuts play a crucial role in determining the complexity of the singularity analysis. In [19], this selection was performed manually, whereas in the present work, the analogous loop-ordering process is fully automated. For example, for the diagram in Fig. 1(a), the algorithm selects the box (with $\mathcal{E}_{1}=3$) as the first loop and the hexagon (with $\mathcal{E}_{2}=5$) as the second. Second, the translation invariance of the integration measure is exploited by shifting the loop momenta, i.e., $\ell_{a}\mapsto\ell_{a}^{\prime}=\ell_{a}+\lambda_{a}$, to eliminate as many scalar products as possible. An explicit example of such shifts is provided in the sunrise example below.
@@
Recently, a semi-automatized Mathematica package BaikovPackage for generating standard and loop-by-loop Baikov representations of Feynman integrals was released [21]. Unlike [Uncaptioned image] , which by default selects the loop-by-loop representation minimizing the number of integration variables $N$, BaikovPackage requires the user to make that choice and specify the order and labeling of the loop momenta $\ell_{a}$. Consequently, the outputs may differ depending on these selections.
@@
$\tilde{\mathcal{G}}(s,m^{2})=s\,,\quad\mathcal{G}(\vec{x},s,m^{2})=\det\left[\begin{array}[]{ccc}m^{2}{+}x_{1}&*&*\\ \frac{1}{2}\left(x_{3}{-}x_{1}{-}x_{2}{-}m^{2}{-}s\right){+}x_{4}{+}x_{5}&m^{2}{+}x_{2}&*\\ x_{4}&x_{5}&s\\ \end{array}\right]\,,$ (2.11)
@@
This makes it clear that only four invariants are needed: $\ell_{1}$, $\ell_{2}^{\prime 2}$, $\ell_{1}\cdot\ell_{2}^{\prime}$, and $p\cdot\ell_{2}^{\prime}$. Taking $x_{4}=\ell_{2}\cdot p$ as the only ISP, the resulting representation has

### v2

@@
Program summary
Program title: [Uncaptioned image] (Singularities of Feynman Integrals Automatized)
Developer’s repository link: https://github.com/StrangeQuark007/SOFIA [repo]
Licensing provisions: MIT license
Programming language: Mathematica 13 or higher
Supplementary material: The example file SOFIA_examples.nb in [repo].
Contents
@@ 1 Introduction
High-precision predictions for particle colliders rely more than ever on the computation of multi-loop Feynman integrals. For this reason, the study of Feynman integrals and their analytic structure has become a major area of research in theoretical particle physics (see, e.g., [Weinzierl:2022eaz, Badger:2023eqz] for reviews). A central challenge in these computations is the reliable prediction of Feynman integral singularities as functions of kinematic parameters, without the need for a direct evaluation of the integral. In this paper, we tackle this problem by introducing [Uncaptioned image] (Singularities of Feynman Integrals Automatized), a Mathematica package designed for this purpose, which we release for public use:
SOFIA GitHub repository [repo].
@@
The most common application is in computing Feynman integrals using the method of differential equations [Kotikov:1990kg]. In the simplest cases, these equations take the form
@@
where $\vec{I}$ is a vector of basis Feynman integrals with $\mathrm{d}$ refers to the differential in the kinematic variables. On the right-hand side, $\mathbf{\Omega}_{i}$ are matrices with entries rational in the kinematics and the spacetime dimension $\mathrm{D}$. In fact, one often looks for bases in which all $\mathbf{\Omega}_{i}$ are linear in the dimensional-regularization parameter $\varepsilon=\frac{4-\mathrm{D}}{2}$ [Henn:2013pwa]. In this case, the resulting Feynman integrals $\vec{I}$ are often expressible in terms of multiple polylogarithms. The arguments of the logarithms, $W_{i}$, are known as the letters [Goncharov:2010jf]. Being able to predict these letters is currently one of the main bottlenecks in computations of polylogarithmic integrals: once $W_{i}$’s are known, one can fit the coefficients $\mathbf{\Omega}_{i}$ and consequently compute $\vec{I}$ using the differential equation together with some boundary conditions.
@@
Figure 1: Examples of cutting‐edge diagrams with timings (in minutes) to obtained singularities and letters. See [repo] for expressions. Dashed lines denote massless particles and otherwise mass labels are displayed along the edges. The choice of SolverBound is determined by the maximum value for which the letters could be subsequently obtained. For diagram (c), although the singularity analysis finished in a reasonable time, the extraction of letters did not complete (hence the “?”) for any reasonable choice of SolverBound. For diagram (a), we used a machine with two Intel Xeon Gold CPUs and 512GB memory, while for (b) and (c) we used a MacBook Pro with an M3 chip and 16GB memory.
@@
Since problems involving analytic structure are notoriously subtle, let us be more precise about the type of questions that we aim to address. A singularity refers to a pole or a branch point in the (complexified) kinematic space. Let us collectively denote the kinematic variables as $\vec{s}\in\mathbb{C}^{\mathfrak{s}}$: they consist of the $\mathfrak{s}$ Mandelstam invariants and masses involved in the problem. While non-perturbatively scattering amplitudes might have a much richer singularity structure, based on examples in perturbation theory [Bogner:2017xhp], it is expected that singularities of the type
@@
where $f_{j}=f_{j}(\vec{s})$ is a polynomial and $A$, $B$ are rational numbers. For example, $\sim\sqrt{f_{j}}\log f_{j}$ is one possibility. Note that this refers to a local behavior of the amplitude near the locus $f_{j}=0$. It is much simpler than the global behavior which has to be compatible with (1.2) for a list of all possible $f_{j}$’s. Indeed, it is known that Feynman integrals are expressed in terms of complicated functions such as multi-polylogarithms, elliptic functions, periods of higher-genus Riemann surfaces and Calabi–Yau manifolds, etc. [Caola:2022ayt, Bourjaily:2022bwx] even though their local behavior (1.2) remains simple. The goal of singularity analysis is to exploit this simplicity.
@@
In the literature, the aforementioned singularities $f_{j}=0$ are also interchangeably known as Landau singularities or anomalous thresholds [Bjorken:1959fd, Landau:1959fi, 10.1143/PTP.22.128, Cutkosky:1960sp]. The equation $f_{j}=0$ is a condition for a network of classical on-shell particles following their classical, though possibly complexified, paths. The simplest case corresponds to positive energies and real angles [Coleman:1965xm], i.e., the physical region. But the interpretation extends to the rest of the physical sheet, and in fact any sheet, after relaxing positivity and reality conditions; see, e.g., [Caron-Huot:2023ikn].
@@
There has been an explosion of recent activity in the topic of determining singularities of Feynman integrals. Let us mention a couple of algorithms that have been implemented as parts of public packages and can handle state-of-the-art Feynman integrals. HyperInt implements a reduction algorithm that finds a superset of Landau singularities [Panzer:2014caa]. PLD.jl implements an algorithm based on toric geometry that finds a computable subset of Landau singularities [Fevola:2023kaw, Fevola:2023fzn]. Both of these approaches work in the Schwinger-like parametrizations of Feynman integrals. A database comparing the two codes can be found in [MathRepo]. The package Baikovletter attempts to find symbol alphabets using Baikov representations [Jiang:2024eaj]. We refer the readers to Sec. 2.2 for more detailed comparison and the introduction of [Fevola:2023fzn] for a more complete list of references.
Importantly, it was recently observed that applying singularity analysis to the elastic unitarity constraint (relating the imaginary part of a diagram to its “cut” sub-diagrams) yields a powerful recursion formula for the calculation of Landau singularities in the loop momentum space [Caron-Huot:2024brh]. This method enables efficient computation of the leading Landau singularities in two-particle reducible diagrams, successfully reproducing known results from the database [MathRepo] while also making new predictions, including all-loop examples such at the families of ladders and penta-ladders.
@@
The first main novelty comes from a new variant of Baikov representations [Baikov:1996iu, Frellesvig:2024ymq] adapted to minimize the number of integration variables and polynomials involved in the singularity analysis. [Uncaptioned image] implements an optimization technique that results in a minimal integral representation (according to a metric explained in Sec. 2.1) without any user input besides the Feynman diagram under consideration. We believe it could become useful in other aspects of Feynman analysis that rely on Baikov representations, such as integration-by-parts reduction or analyzing Feynman geometries [Frellesvig:2017aai, Bogner:2019lfa, Frellesvig:2019uqt, Chen:2022lzr, Giroux:2022wav, Pogel:2022vat, Delto:2023kqv, Giroux:2024yxu, Duhr:2024bzt, Frellesvig:2024zph]. We illustrate the latter application in Sec. 4.2. Likewise, we extend our analysis to eikonalized Feynman integrals, such as those in post-Minkowskianan expansions with an example in Sec. 4.3.
@@
The second key innovation is the integration of advanced polynomial system solving methods for singularity analysis into the optimized Baikov representation. In particular, we implement a variant of the Fubini reduction algorithm [Brown:2009ta] that results in an upper bound on the singularity locus. We also implement a Mathematica interface to PLD.jl, which allows one to seamlessly perform a numerical and symbolic analysis on a given diagram. We emphasize that this algorithm can be applied to any Euler integral [Matsubara-Heo:2023ylc] and give a couple of examples of applications to cosmological and energy correlators in Sec. 4.4.
@@
The third and final significant novelty is the implementation of graph‐theoretical tools to break down and speed up the computation of singularities. It has long been observed that the symmetries of a Feynman diagram lead to an identification between Schwinger parameters of the leading physical sheet singularity [kolkunov1960singular, KORS] (see [Correia:2021etg] for a modern application in the context of the S-matrix bootstrap). In this work, we extend this idea by relating different sets of subleading singularities through the construction and identification of subtopologies that are connected via symmetry transformations accounting for variations in their kinematics. This procedure effectively mitigates the factorial growth in the number of subtopologies that would otherwise need to be analyzed.
@@
Finally, we use the recent implementation of the publicly available packge Effortless [Effortlessxxx] to provide an end-to-end function that takes a diagram as the input and provides a candidate symbol alphabet as the output. By “candidate” we mean that the alphabet might be overcomplete. An example is provided in Sec. 4.1.
@@
We believe that [Uncaptioned image] establishes substantial new milestones in the singularity analysis of Feynman integrals, producing results that were previously unattainable. The ancillary files [repo] contain several new predictions for singularities and alphabets, including those presented in Fig. 1, which exemplify how [Uncaptioned image] can be applied to next-frontier examples, namely, $HZ+2$ jets, H+H+jet, and multi-jets production through light- (for (a)) or heavy-quark (for (b) and (c)) loop‐induced gluon fusion at the LHC.
@@
We hope that [Uncaptioned image] will serve as a tool to explore new avenues and sheds light on the analytic structures of perturbative scattering amplitudes that were previously inaccessible.
@@
We provide theoretical background in Sec. 2, including the discussion of different variants of the Baikov representation, singularity analysis of integrals, and mapping from singularities to symbol alphabets. The manual of [Uncaptioned image] is given in Sec. 3 with an outline of all the functions and options. Sec. 4 focuses on examples and various applications. Finally, future directions are discussed in Sec. 5.
@@ 2.1 Baikov representations
There are a few ways to implement the change of variables (2.3). The general idea dates back to the S-matrix theory in the 1960s [Mandelstam:1959bc, Gribov:1962ft, Correia:2020xtr], although in the more recent literature on Feynman integrals parametrizations of this type are referred to as Baikov representations [Baikov:1996iu]. In the global (also called standard or democratic or original) Baikov representation, one transforms all variables at the same time. This results in the representation of the form
@@
The central object in the global Baikov representation is $\mathcal{G}(\vec{x},\vec{s})$ which arises as the general Jacobian for the change of variables. It can be expressed as a Gram determinant of a certain matrix. The integration contour $\Gamma^{\prime}$ is determined by $\vec{x}$ such that $\mathcal{G}(\vec{x},\vec{s})>0$. Likewise, $\tilde{\mathcal{G}}(\vec{s})$ is another Gram determinant that depends only on external kinematics and hence acts as an overall normalization. The derivation in App. A makes it clear that one can include different powers of propagators as well as non-trivial numerators with minimal modifications. For completeness, we review these aspects in App. A (see also [Lee:2009dh, Mastrolia:2018uzb]).
The alternatives proposed by Frellesvig and Papadopoulos are loop-by-loop Baikov representations [Frellesvig:2017aai]. The idea is simply to Baikov parametrize each loop of the diagram sequentially in a given order. For each loop, we look for a shift in the loop momentum $\ell_{a}$ so that fewer invariants—such as $\ell_{a}\cdot p_{i}$ and $\ell_{a}\cdot\ell_{b}$—appear before changing variables from the $\ell$’s to the $x$’s. This procedure has the effect of decreasing the total number of integration variables $N\leqslant N^{\prime}$. Typically, the minimal/optimal $N$, is [Frellesvig:2024ymq]
@@
Due to complicated integration contours, Baikov representations are typically not used in the explicit evaluation of Feynman integrals. However, they are extremely useful for integral reduction, integration-by-parts identities, singularity analysis, and other applications that rely on the simplicity of the integrand. The main advantage comes from the application of generalized unitarity. Generalized cuts can be taken by simply replacing the integration contour with residues around the relevant propagators (similarly, unitarity or Cutkosky cuts can be taken after an extra complex conjugation). For example, the maximal cut is obtained by
@@
in either the global or loop-by-loop Baikov approach (i.e., $\mathcal{N}=N\,,N^{\prime}$). This replacement exists if the original integration domain, $\Gamma$ or $\Gamma^{\prime}$, restricted to $x_{1}=x_{2}=\ldots=x_{\mathcal{N}}=0$ is non-empty (more on that is discussed in Sec. 3.3). Note that after taking the cut, $\mathcal{N}-E$ variables still have to be integrated over.
@@
We implement a search algorithm that optimizes for the most efficient loop-by-loop Baikov representation without any user input. It proceeds in two steps. First, when possible, the loops are ordered automatically so that the number of external momenta $\mathcal{E}_{a}$ (associated with each loop momentum $\ell_{a}$, for $1\leqslant a\leqslant L$) increases from the smallest to the largest. [^1] This strategy is inspired by the recursive elastic unitarity approach of [Caron-Huot:2024brh], where the selection and ordering of cuts play a crucial role in determining the complexity of the singularity analysis. In [Caron-Huot:2024brh], this selection was performed manually, whereas in the present work, the analogous loop-ordering process is fully automated. For example, for the diagram in Fig. 1(a), the algorithm selects the box (with $\mathcal{E}_{1}=3$) as the first loop and the hexagon (with $\mathcal{E}_{2}=5$) as the second. Second, the translation invariance of the integration measure is exploited by shifting the loop momenta, i.e., $\ell_{a}\mapsto\ell_{a}^{\prime}=\ell_{a}+\lambda_{a}$, to eliminate as many scalar products as possible. An explicit example of such shifts is provided in the sunrise example below.
@@
Recently, a semi-automatized Mathematica package BaikovPackage for generating standard and loop-by-loop Baikov representations of Feynman integrals was released [Frellesvig:2024ymq]. Unlike [Uncaptioned image] , which by default selects the loop-by-loop representation minimizing the number of integration variables $N$, BaikovPackage requires the user to make that choice and specify the order and labeling of the loop momenta $\ell_{a}$. Consequently, the outputs may differ depending on these selections.
@@
$\hskip-8.5359pt\tilde{\mathcal{G}}(s,m^{2})=s\,,\quad\mathcal{G}(\vec{x},s,m^{2})=\det\left[\begin{array}[]{ccc}m^{2}{+}x_{1}&*&*\\ \frac{1}{2}\left(x_{3}{-}x_{1}{-}x_{2}{-}m^{2}{-}s\right){+}x_{4}{+}x_{5}&m^{2}{+}x_{2}&*\\ x_{4}&x_{5}&s\\ \end{array}\right]\,$, (2.11)
@@
This makes it clear that only four invariants are needed: $\ell_{1}^{2}$, $\ell_{2}^{\prime 2}$, $\ell_{1}\cdot\ell_{2}^{\prime}$, and $p\cdot\ell_{2}^{\prime}$. Taking $x_{4}=\ell_{2}\cdot p$ as the only ISP, the resulting representation has